OpenStudy (anonymous):

please help!! 1 medal and fan OR testimony Read the following two statements. Then, if possible, use the Law of Detachment to draw a conclusion. If x = 3, then 3x – 4 = 5. x = 3 x = 3 If 3x – 4 = 5, then x = 3. 3x – 4 = 5 not possible

OpenStudy (anonymous):

If 3x – 4 = 5, then x = 3....

OpenStudy (anonymous):

why? explain

OpenStudy (anonymous):

If 3x – 4 = 5 3x=5+4=9 x=9/3 x=3

OpenStudy (anonymous):

good